Using (x + a) (x + b) = x^2 + (a + b) x + ab, find
9.7 x 9.8

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

9.7 x 9.8
= (9 + 0.7 )(9 + 0.8)
= 9^2+ (0.7 + 0.8)9 + 0.56
= 81 + 13.5 + 0.56
= 95.06

Thanks